of the 30 students in a class, 13 love swimming, 15 love paddle boarding and 7 like neither.
a) how many students love both swimming and paddle boarding
b) how many love swimming but not paddle boarding
c) what is the probability of picking a student who loves paddle boarding, but not

Answer by ikleyn(53763)

.
of the 30 students in a class, 13 love swimming, 15 love paddle boarding and 7 like neither.
a) how many students love both swimming and paddle boarding
b) how many love swimming but not paddle boarding
c) what is the probability of picking a student who loves paddle boarding, but not
~~~~~~~~~~~~~~~

Since of the 30 students 7 like neither, it means that 30-7 = 23 like something (one of the two).


Therefore,


a)  how many students love both swimming and paddle boarding = (13+15) - 23 = 28-23 = 5.


b)  how many love swimming but not paddle boarding = 13 - 5 = 8.


c)  I can not answer c) because this question is uncompleted in your post.

-------------
See the lesson
    - Counting elements in sub-sets of a given finite set
in this site.